Hi! I need some help solving this equation and if possible expressing my answer as on ordered pair.
y=2x+1
x+y=4
Thanks in advance for your help!

Answer by tutorcecilia(2152)

y=2x+1
x+y=4 [Pick either equation and solve for one of the variables]
x+y=4
y=-x+4 [Plug the value of (y=-x+4) into the first equation and solve for x]
.
y=2x+1
-x+4=2x+1 [solve for x]
4=3x+1
3=3x
x=1 [Plug (x=1) back into either equation and solve for y]
.
x+y=4
1+y=4 [Solve for y]
y=3
.
Check by plugging the values of x and y back into either equation.
y=2x+1
3=2(1)+1
3=2+1
3=3 [Checks out]